# Bayesian Group Regularization in Generalized Linear Models with a Continuous Spike-and-Slab Prior thanks: Keywords and phrases: generalized linear model, group sparsity, spike-and-slab, spike-and-slab group lasso, posterior contraction, rate of convergence, variable selection

### v8

@@ B.1 Proof of Theorem 2
Similar to (D.24) in the proof of Theorem 2 of Bai et al. [3], the first term of (B.6) can be shown to be the dominating term, and thus, $-\log\Pi(\mathcal{B}_{n})\lesssim n\epsilon_{n}^{2}$. This establishes (B.2) and the proof is finished. $\square$
@@
[…]
@@
[…]
@@ B.2 Proof of Theorem 3
[…]
@@
[…]
@@
where we used the tail bound for the gamma density on p. 29 of Boucheron et al. [8] and the inequality $21+x-\sqrt{1+2x}\geq(x-1)/$ 2 for $x>0$ (similarly as in Ning et al. [31]) to bound the integral in the second line from above by $\exp(-\lambda_{0}\omega_{g}+m_{\max})$.
@@
(${G\choose K_{1}s_{0}}D\left(\frac{C_{5}\epsilon_{n}}{36G\sqrt{2}},\left\{\lVert\bm{\beta}_{S}-\bm{\beta}_{0S}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\},\lVert\cdot\rVert_{2}\right)\leq{G\choose K_{1}s_{0}}\left(\frac{108\sqrt{2}G^{{\color[rgb]{0,0,0}M}+1}}{C_{5}\epsilon_{n}}\right)^{K_{1}s_{0}},$ (B.17)
@@
where we used the fact that (${G\choose K_{1}s_{0}}\leq G^{K_{1}s_{0}}$. Having established (B.18), we can now use Lemma 9 of Ghosal and van der Vaart [17], which implies that for every $\epsilon>\epsilon_{n}$, there exists a test $\bar{\varphi}_{n}$ such that
@@
where we can choose $C_{4}>C_{3}$ so that $-\lambda_{1}G^{{\color[rgb]{0,0,0}M}}$ is the dominating term in the first line of the display, and further, $G^{{\color[rgb]{0,0,0}M}}\gg C_{4}s_{0}n\log G$ when $n$ is large. Thus, from (B.23), we can see that (B.19) holds, and we have established the posterior contraction rate with respect to the average squared Hellinger metric (B.13).
@@ B.3 Proof of Proposition 1
[…]
@@ C.2 M-step for updating the regression coefficients
where $-\ell_{n}(\bm{\beta})$ is the negative of the log-likelihood in (1.4). While (C.3) may appear to be intractable, we can apply the standard iteratively reweighted least squares (IRLS) algorithm [28] for GLMs to efficiently solve (C.3).
@@
$-\ell_{n}(\bm{\beta})\approx\frac{1}{2}(\mathbf{Z}-\mathbf{X}\bm{\beta})^{\top}\mathbf{W}(\mathbf{Z}-\mathbf{X}\bm{\beta}).$ (C.4)
Thus, substituting the approximation (C.4) for $-\ell_{n}(\bm{\beta})$ in (C.3), we see that the M-step (C.3) for $\bm{\beta}$ can alternatively be written as

### v9

@@ B.1 Proof of Theorem 2
Similar to (D.24) in the proof of Theorem 2 of Bai et al. [3], the first term of (B.6) can be shown to be the dominating term, and thus, $\log\Pi(\mathcal{B}_{n})\lesssim n\epsilon_{n}^{2}$. This establishes (B.2) and the proof is finished. $\square$
@@
$\mathbb{E}_{0}\left(\int_{\mathcal{A}_{n}^{c}}\prod_{i=1}^{n}\frac{f_{i}(y_{i})}{f_{0i}(y_{i})}d\Pi(\bm{\beta})\right)\leq\int_{\mathcal{A}_{n}^{c}}d\Pi(\bm{\beta})=\Pi(|\bm{\beta}|>K_{1}s_{0}).$ (B.8)
@@
$\Pi(|\nu(\bm{\beta})|>K_{1}s_{0})$ $\leq\sum_{S:|S|>C_{3}s_{0}}2^{|S|}\theta^{|S|}\left\{\prod_{g\in S}\int_{\lVert\bm{\beta}_{g}\rVert_{2}>\omega_{g}}C_{g}\lambda_{1}^{m_{g}}e^{-\lambda_{1}\lVert\bm{\beta}_{g}\rVert_{2}}d\bm{\beta}_{g}\right\}\times\left\{\prod_{g\in S^{c}}\int_{\lVert\bm{\beta}_{g}\rVert_{2}\leq w_{g}}\pi(\bm{\beta}_{g})d\bm{\beta}_{g}\right\}$
@@ B.2 Proof of Theorem 3
$\Pi(\lVert\bm{\beta}_{g}\rVert_{2}>\omega_{g})$ $\geq\int_{\omega_{g}}^{\infty}\frac{1}{\Gamma(m_{g})\lambda_{0}^{m_{g}}}x^{m_{g}-1}e^{-x/\lambda_{0}}dx$
$\geq\frac{1}{\Gamma(m_{g})\lambda_{0}^{m_{g}}}\int_{\lambda_{0}}^{\infty}x^{m_{g}-1}e^{-x/\lambda_{0}}dx$
$\geq\frac{1}{\Gamma(m_{g})\lambda_{0}}\int_{\lambda_{0}}^{\infty}e^{-x/\lambda_{0}}dx$
@@
$\Pi(\lVert\bm{\beta}_{g}\rVert_{2}\leq\omega_{g})$ $\geq\int_{0}^{\omega_{g}}\frac{1}{\Gamma(m_{g})\lambda_{0}^{m_{g}}}x^{m_{g}-1}e^{-x/\lambda_{0}}dx$
$=1-\int_{\omega_{g}}^{\infty}\frac{1}{\Gamma(m_{g})\lambda_{0}^{m_{g}}}x^{m_{g}-1}e^{-x/\lambda_{0}}dx$
@@
where we used the tail bound for the gamma density on p. 29 of Boucheron et al. [8] and the inequality $1+x-\sqrt{1+2x}\geq(x-1)/2$ for $x>0$ (similarly as in Ning et al. [31]) to bound the integral in the second line from above by $\exp(-\lambda_{0}\omega_{g}+m_{\max})$.
@@
${G\choose K_{1}s_{0}}D\left(\frac{C_{5}\epsilon_{n}}{36G\sqrt{2}},\left\{\lVert\bm{\beta}_{S}-\bm{\beta}_{0S}\rVert_{2}\leq G^{{\color[rgb]{0,0,0}M}}\right\},\lVert\cdot\rVert_{2}\right)\leq{G\choose K_{1}s_{0}}\left(\frac{108\sqrt{2}G^{{\color[rgb]{0,0,0}M}+1}}{C_{5}\epsilon_{n}}\right)^{K_{1}s_{0}},$ (B.17)
@@
where we used the fact that ${G\choose K_{1}s_{0}}\leq G^{K_{1}s_{0}}$. Having established (B.18), we can now use Lemma 9 of Ghosal and van der Vaart [17], which implies that for every $\epsilon>\epsilon_{n}$, there exists a test $\bar{\varphi}_{n}$ such that
@@
where we can choose $C_{4}>C_{3}$ so that $\lambda_{1}G^{{\color[rgb]{0,0,0}M}}$ is the dominating term in the first line of the display, and further, $G^{{\color[rgb]{0,0,0}M}}\gg C_{4}s_{0}n\log G$ when $n$ is large. Thus, from (B.23), we can see that (B.19) holds, and we have established the posterior contraction rate with respect to the average squared Hellinger metric (B.13).
@@ B.3 Proof of Proposition 1
Since $\mathbf{Y}_{g}\sim\mathcal{N}_{m}(\bm{0}_{m},\mathbf{I}_{m})$ for all $g\neq 1$, $\lVert\mathbf{Y}_{g}\rVert_{2}^{2}\sim\chi_{m}^{2}$. Letting $\Gamma(a,x)=\int_{x}^{\infty}t^{a-1}e^{-t}dt$ denote the incomplete gamma function, we proceed to upper-bound the individual summands as
$\mathbb{E}_{0}\lVert\widehat{\bm{\beta}}_{g}\rVert_{2}^{2}$ $\leq\int_{\lambda^{2}}^{\infty}\left[\Gamma\left(\frac{m}{2}\right)\right]^{-1}\left(\frac{u}{2}\right)^{m/2}e^{-u/2}du$
@@ C.2 M-step for updating the regression coefficients
where $\ell_{n}(\bm{\beta})$ is the negative of the log-likelihood in (1.4). While (C.3) may appear to be intractable, we can apply the standard iteratively reweighted least squares (IRLS) algorithm [28] for GLMs to efficiently solve (C.3).
@@
$\ell_{n}(\bm{\beta})\approx\frac{1}{2}(\mathbf{Z}-\mathbf{X}\bm{\beta})^{\top}\mathbf{W}(\mathbf{Z}-\mathbf{X}\bm{\beta}).$ (C.4)
Thus, substituting the approximation (C.4) for $\ell_{n}(\bm{\beta})$ in (C.3), we see that the M-step (C.3) for $\bm{\beta}$ can alternatively be written as